OpenStudy (anonymous):

solve the equation b+2/3 - b-2/7 =2 5.1 5.5 6.0 6.5 explain please

OpenStudy (anonymous):

i assume this is \[\frac{b+2}{3}-\frac{b-2}{7}=2\]

OpenStudy (anonymous):

your best bet is to multiply both sides by the least common multiple of \(3\) and \(7\), namely \(21\)

OpenStudy (anonymous):

\[21\left(\frac{b+2}{3}-\frac{b-2}{7}\right)=2\times 21\]or \[7(b+2)-3(b-2)=42\] then solve that one

OpenStudy (anonymous):

14b-1b